OpenStudy (anonymous):

A children's birthday party at an indoor play center costs $50 to rent the space and $6.25 per child. Gina wants to spend no more than $100 on her son's fourth birthday party. Write an inequality for the number of children who can attend the birthday party, and solve the inequality. A) 50 + 6.25x ≤ 100; x ≤ 8 B) 50 + 6.25x ≥ 100; x ≥ 8 C) 50 + 6.25x < 100; x < 8 D) 50 + 6.25x > 100; x > 8 @gerryliyana

OpenStudy (anonymous):

It's A 50 + 6.25x \[50 + 6.25x \le 100; x \le 8\]

OpenStudy (anonymous):

\[50 + 6.25x \le 100; x \le8\]

OpenStudy (anonymous):

Ok so what do you have to do to get that

OpenStudy (anonymous):

number of childrens = x total cost <= 100 cost for x children's = 6.25x so final equation should be 6.25x + 50 <= 100

OpenStudy (anonymous):

Oh ok thank you! I have some more questions.

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

Ok next question: The gray area is the sidewalk. The area of the sidewalk is ___________ square units. A) 48 B) 56 C) 72 D) 84 Pic: http://www.usatestprep.com/modules/gallery/files/4/409/409.gif

OpenStudy (anonymous):

The area of the sidewalk = the gray box area - the green box are have you tried it?

OpenStudy (anonymous):

I sorta don't know how to do it.

OpenStudy (anonymous):

first, you are supposed to find the gray (big) area (including green area in it) and the green area

OpenStudy (anonymous):

Ok so that would be 120

OpenStudy (anonymous):

yes, and the green area would be 48

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So would we divide that?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Oh so what do we do next?

OpenStudy (anonymous):

The area of the sidewalk = the gray box area - the green box area

OpenStudy (anonymous):

The area of the sidewalk = the gray box area - the green box are The area of the sidewalk = 120 - 48

OpenStudy (anonymous):

And that would be 72

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

So it would be 72

OpenStudy (anonymous):

yes good job!

OpenStudy (anonymous):

the answer is C

OpenStudy (anonymous):

Awesome! Well that's it for me tonight. Thank you so much! Will you be on Open Study tomorrow?

OpenStudy (anonymous):

you are very welcome. Yes I will

OpenStudy (anonymous):

Ok I will see you tomorrow then! Bye!

OpenStudy (anonymous):

ok, bye